OpenStudy (anonymous):

Given the equation y − 3 = 1/2(x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. x − 2y = −12 y = 1/2x y = 1/2x + 6 y = 1/2x + 9

jimthompson5910 (jim_thompson5910):

were you able to get anywhere?

OpenStudy (anonymous):

no.

jimthompson5910 (jim_thompson5910):

ok sorry for the delay

jimthompson5910 (jim_thompson5910):

the idea is to get the equation into y = mx+b form

jimthompson5910 (jim_thompson5910):

so all we need to do is isolate y and get the equation into that form like this y − 3 = 1/2(x + 6) y = 1/2(x + 6) + 3 y = 1/2x + (1/2)*6 + 3 y = 1/2x + 3 + 3 y = 1/2x + 6 hopefully all this makes sense If not, then go over it a few times and hopefully a few more things will click

OpenStudy (anonymous):

It's fine:)

OpenStudy (anonymous):

Yes...Thanks I am terrible at math.

jimthompson5910 (jim_thompson5910):

keep practicing and you'll get better

OpenStudy (anonymous):

Thanks so much.

jimthompson5910 (jim_thompson5910):

you're welcome